OpenStudy (anonymous):

Simplify 4 x 5 – 4 2 x 6

OpenStudy (unklerhaukus):

is anything between the 4 and the 2?

OpenStudy (anonymous):

is there suppose to be a symbol between the 4 and the 2?

OpenStudy (anonymous):

yes sorry a division sign

OpenStudy (unklerhaukus):

\[4 \times 5 – 4\div2 \times 6 \]

OpenStudy (anonymous):

correct

OpenStudy (unklerhaukus):

\[(4 \times 5) –( (4\div2) \times 6)\]

OpenStudy (anonymous):

then it wuld be 8

OpenStudy (anonymous):

Thanks:) i'm still gonna try to work it out so i get it

OpenStudy (anonymous):

np